# LetsTalk: Latent Diffusion Transformer for Talking Video Synthesis

### v1

@@ 6.1 Backbone
Table 5: Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
6.2 Audio Encoder
@@
We resample the audio signal to a 16 KHz sample rate. For audio feature extraction, we use wav2vec [36], employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation. To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips. Additionally, we use a sliding window (multiple audio frames) for each video frame to maintain audio continuity and preserve more accurate audio information. To extract audio representations suitable for transformer tokenization, we designed an audio projection module. This module first flattens the extracted audio features and then selects a fixed number of context tokens (32 in our experiments). Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
6.3 Siamese Transformer
@@
The Siamese Fusion scheme has been adopted in AnimateAnyone [21], EMO [41], Hallo [47], and so on. The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise. Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network. Considering LetsTalk as a Diffusion Transformer, we introduce a ViT [14] pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image. For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
7 Additional Implementation Details
7.1 Diffusion
@@
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM [13], specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
7.2 Long-Duration Generation
Figure 8: The illustration of the long-duration generation process.
Unlike the previous frame-by-frame generation approach [38, 39], our end-to-end diffusion transformer generates a video clip in a single sampling process. In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences. Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible. To address this, we follow [47] and pick the last $N$ frames from the last generated clip as the initial frames of the current clip ($N=2$ in our experiments). The initial frames are solely involved in guiding the temporal attention, ensuring continuity between consecutive clips during sampling. Fig. 8 illustrates the long-duration generation process.
7.3 Training
@@
We initialize LetsTalk from a pre-trained Latte [29] model on the FaceForensics dataset [35]. Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion. For training, we use the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$ and no weight decay. The batch size ranges from 10 to 16, and the only data augmentation applied is horizontal flips. Each model is trained for 300k to 500k iterations. Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
7.4 Inference
@@
8 Qualitative Comparisons
In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk. We show more results on the HDTF [53] and Celeb-HQ [56] benchmarks ( Fig. 9 and Fig. 10).
@@ 8 Qualitative Comparisons
Figure 9: The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
Figure 10: The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
9 Visualization
We provide video visualizations in our attachment. Notably, the model used to generate the video is only trained on HDTF [53] and CelebV-HQ [56], which only include English conversation scenarios of real portraits. In addition, we use Chinese audio and singing audio to drive real portraits. Furthermore, LetsTalk can drive AI-generated portraits [^6] We use Doubao to generate AI-generated portraits in the attached video.. The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
Figure 11: English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 12: Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 13: Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 14: Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.
10 Limitation and Future Work
Despite the advancements proposed in this study regarding portrait image animation, several limitations necessitate further exploration and consideration. (1) Resource and Dataset Limitations: Unlike recent works such as Hallo [47] and EMO [41], which use diverse real-world datasets, our study is limited by the availability of publicly accessible datasets. These datasets are relatively small and may not be as robust, thus limiting the generalizability of LetsTalk. Additionally, the computational resources available were insufficient to fully exploit larger models, restricting us to smaller-scale models with fewer parameters. Future research could benefit from larger datasets and scaling up model sizes to explore the scalability of LetsTalk. (2) Long-Duration Animation: A key limitation in generating long-duration animations with our end-to-end framework is the challenge of maintaining smooth transitions between video clips. Due to the high realism and dynamism of the generated clips, seamless continuity is difficult to achieve. While the method in  Fig. 8 mitigates this by using motion frames, the discrepancy between real frames used during training and synthetic frames used during inference leads to cumulative errors. As the video length increases, these errors accumulate, resulting in a degradation of quality in the later clips. Future work will focus on minimizing this error propagation, possibly through improved temporal modeling or alternative approaches to frame synthesis during inference. (3) High-Resolution: Although Latent Diffusion Models (LDM) [34] leverage VAE to enhance the resolution of synthesized images, high resolution remains a significant challenge for video tasks. Recent works in portrait animation, such as [41, 45, 47], have circumvented this issue through a two-stage training process. However, we hope that future research will explore more effective solutions for generating high-resolution outputs in end-to-end models like LetsTalk.

### v2

@@ Appendix A Preliminaries
Table 7. Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
Appendix B Network Architecture Details
B.1. Backbone
In LetsTalk, we borrow the Variational AutoEncoder (VAE) (Kingma, 2013) from Stable Diffusion (Rombach et al., 2022). Given the reference portrait image and noisy video, we first encode them using the encoder in VAE. Then we patchify and flatten them along the spatial dimension. Finally, we use an MLP to project them into the latent space. The embeddings in the latent space will be fed into our transformer backbone. Specifically, our LetsTalk model is constructed using a series of N transformer blocks, each with a hidden dimension of D and H multi-head attention heads. Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab. 7. Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$. Next, the two features are partitioned according to $2\times 2$ patch size and flattened as $\bm{P}\in\mathbb{R}^{256\times 16}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 16}$. Take our base model (i.e. LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
B.2. Adaptive Integration
Following VDT(Lu et al., 2023), we insert an adaptive full-connected layer between attention layers:
(4) $\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$
where $\alpha$ is a learnable scalar and $\text{FC}(\cdot)$ denotes a fully-connected layer. This enables adaptive fusion of temporal dynamics into spatial representations.
B.3. Audio Encoder
@@
We resample the audio signal to a 16 KHz sample rate. For audio feature extraction, we use wav2vec (Schneider et al., 2019), employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation. To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips. Additionally, we use a sliding window (multiple audio frames) for each video frame to maintain audio continuity and preserve more accurate audio information. To extract audio representations suitable for transformer tokenization, we designed an audio projection module. This module first flattens the extracted audio features and then selects a fixed number of context tokens (32 in our experiments). Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
B.4. Siamese Transformer
@@
The Siamese Fusion scheme has been adopted in AnimateAnyone (Hu, 2024), EMO (Tian et al., 2024), Hallo (Xu et al., 2024b), and so on. The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise. Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network. Considering LetsTalk as a Diffusion Transformer, we introduce a ViT (Dosovitskiy, 2020) pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image. For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
Figure 7. Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
B.5. Impact of Audio-guidance on Diversity
To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 00 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 7 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
Appendix C Additional Implementation Details
C.1. Diffusion
@@
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
C.2. Training
@@
We initialize LetsTalk from a pre-trained Latte (Ma et al., 2024) model on the FaceForensics dataset (Rössler et al., 2018). Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion. For training, we use the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$ and no weight decay. The batch size ranges from 10 to 16, and the only data augmentation applied is horizontal flips. Each model is trained for 300k to 500k iterations. Following common practices in generative modeling works (Peebles and Xie, 2023; Bao et al., 2023), an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model. Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training. Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
C.3. Inference
@@
C.4. Diffusion
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
Appendix D Qualitative Comparisons
In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk. We show more results on the HDTF (Zhang et al., 2021) and Celeb-HQ (Zhu et al., 2022) benchmarks ( Fig. 8 and Fig. 9).
@@ Appendix D Qualitative Comparisons
Figure 8. The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
Figure 9. The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
Appendix E Visualization
We provide video visualizations in our attachment. Notably, the model used to generate the video is only trained on HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022), which only include English conversation scenarios of real portraits. In addition, we use Chinese audio and singing audio to drive real portraits. Furthermore, LetsTalk can drive AI-generated portraits [^2] We use Doubao to generate AI-generated portraits in the attached video.. The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
Figure 10. English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 11. Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 12. Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 13. Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.